OpenStudy (anonymous):

what is the smallest total resistance which can be obtained using a 6 ohm and a 12 ohm resistor. 2ohm , 4ohm, 6ohm,8ohm

OpenStudy (anonymous):

1/Req = 1/R1+1/R2|| 1/Req= (R1+R2)/(R1*R2) || Req = (R1*R2)/ (R1+R2) || Req=(6*12)/(6+12)=4ohm .........the second choice is right...

OpenStudy (anonymous):

how many more have you?